Required information [The following information applies to the questions displayed below.) Part 1 of 3 0.62 points BMX CompanX Answer is complete but not entirely correct. a) Tax Тах Tax Rate Amount 6.20% $ 99.20 FICA—Social Security FICA-Medicare SeAssuming situation (a), prepare the employers September 30 journal entries to record the employers payroll taxes expense anAssuming situation (a), prepare the employers September 30 journal entries to record the employers payroll taxes expense an
Answer #1
1
a)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 1600 6.20% 99.20
FICA-Medicare 1600 1.45% 23.20
FUTA 800 0.60% 4.80
SUTA 800 5.40% 43.20
b)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 2950 6.20% 182.90
FICA-Medicare 2950 1.45% 42.78
FUTA 2950 0.60% 17.70
SUTA 2950 5.40% 159.30
c)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 4600 6.20% 285.20
FICA-Medicare 9700 1.45% 140.65
FUTA 0 0.60% 0.00
SUTA 0 5.40% 0.00
Note: Earnings taxable for FUTA and SUTA in (a) = 7000-6200 = $800
Note: Earnings taxable for Social Security in (c) = 128400-123800 = $4600
2
Employer Payroll taxes September earnings subject to tax Tax Rate Tax Amount
FICA-Social Security 1600 6.20% 99.20
FICA-Medicare 1600 1.45% 23.20
FUTA 800 0.60% 4.80
SUTA 800 2.90% 23.20
Total Payroll Taxes Expense 150.40
Date General Journal Debit Credit
Sep. 30 Payroll taxes expense 150.40
FICA—Social security taxes payable 99.20
FICA—Medicare taxes payable 23.20
Federal unemployment taxes payable 4.80
State unemployment taxes payable 23.20